MODS1 Commissioning Run 3: UTC 2010 December 17

Observer(s): R. Pogge and C. Henderson (OSU)
Telescope Operator: D. Gonzalez (LBTO)
Telescope Support: J. Urban (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Cloud deck above with occasional wisps around us until 2pm MST, then we were enshrouded by fog. Light snow began at 3pm MST. Humidity 100% since last night, and temperatures steady at -3C most of the afternoon. Snow through morning in the forecast.

A UPS on level 1 went out about 9am taking out all of the computer rooms and network until about noon. MODS startup had one glitch we are investigating but otherwise we came up fine and are ready to go when the sky is. We'll be doing focus tests, calibrations, and systems check out tonight.

Worked through the various types of internal calibration data for all dual-channel mods (imaging, grating, and prism), biases, flats, and comps.

Snowed most of the first part of the night, clearing a little by dawn twilight. Temp at twilight -6C, 100% humidity, wind 10-15m.s out of the SW.

Preparations

Dewars filled at 1615 MST, OK as before.

Minor modifications to lbttcs agent following suggestions from LBTO personnel for DD entries to use for parallactic and position angle parameters.

Modifications made to the MODS control panel to implement changes and upgrades derived from the earlier runs. Tested new feature that will automatically turn off the calibration lamps after a preset time if left unattended (lamps already turn off automatically when the calibration tower is retracted). The lamp timeout is global for all lamps, and can be set either in the runtime config file (sets the default) or by an engineering command. Current default is 3600s (1h).

Details

All Times are UTC unless otherwise noted.

0015 - snow increasing in intensity with near white-out conditions. Snow is failry dry and icy. Wind gusts near 20m/s, humidity 100% temp -3C and falling slowly.

0340 - After LUCI flexure tests, starting to take biases and calibrations with the dome lights turned off. Weather still poor with a winter weather advisory in effect for the upper elevations including Mt. Graham.

Bias Frames

Biases
  • Full-frame 8Kx3K Spectroscopic unbinned x10 in each channel.
    Script: bias8K.cal
    Execution Time: 15 minutes
  • 3Kx3K Imaging unbinned x10 in each channel.
    Script: bias3K.cal
    Execution Time: 9 minutes
    Note: lights on in the dome during the first few images

0440 - Pause for David working on UMAC problem with MODS authorization. Proceeding anyway

Grating Mode Calibrations

Grating Spectroscopy, dual mode calibrations.
  • Slit: LS5x60x0.6
  • Grating: Blue=G400L Red=G670L
  • CenLam: Blue=4350A (nominal) Red=7730A (nominal) - will measure

Comp Lamps - Script: grlamps.cal
  • IMCS Center and Lock
    Execution Time: 45s
  • Ne+Hg(Ar) Lamps x 3
    Red: ND1.5 + 1s
    Blue: NC1.5+2s
    30s for Hg(Ar) lamp to warm up
    Execution Time: 5min
  • Hg(Ar) Lamp x 3
    Red: Idle
    Blue: UG5 + 2s
    Execution Time: 4min
    Could pair with Ne or Ar in red at no extra time cost
  • Kr+Xe Lamps x 3
    Red: ND1.5 + 1s
    Blue: Clear + 15s
    Execution Time: 4min
  • Ar Lamp x 3
    Red: ND1.5 + 1s
    Blue: Clear + 30s
    Execution Time: 5min
    Some ~45sec of deadtime after completion of red images

Total execution time for the sequence including IMCS Lock: ~20 minutes. Does not include time to configure from imaging mode, approximately 2 minutes including mask insert.

0550 - Spectral Flat Fields - Script: grflats.cal

  • IMCS Center and Lock
    Execution Time: ~20 sec
  • QTH1 Lamp x 10
    Red: ND1.5 + 2s
    Blue: Clear + 1s
    Execution Time: 15min
  • QTH1+2 Lamps x 10
    Red: ND1.5 + 1s
    Blue: UG5 + 6s
    Execution Time: 15min
  • VFLAT 10.0 x 10
    Red: Clear + 4s
    Blue: Idle
    Execution Time: 15min
    Lamp too red for useful blue flats

Total execution time is about 45 minutes

0640 - Still snowing, wind now 10-15m/s, Temp -4.3C, Humidity 100%

Imaging Mode Calibrations

0730 - Red Camera Imaging Flats - Script: redimflat.cal

Using the VFLAT lamp. Not possible to exposure balance readily between blue and red.

  • IMCS Center and Lock
    Execution Time: ~20 sec
  • SDSS r: VFLAT 2.0 + 25s x 10
    Execution Time: 12min
  • SDSS i: VFLAT 1.5 + 15s x 10
    Execution Time: 11 min
  • SDSS z: VFLAT 1.5 + 8s x 10
    Execution Time: 10 min

Total execution time is about 35 min with setup.

0815 - Blue Camera Imaging Flat - Script: blueimflat.cal

  • IMCS Center and Lock
    Execution Time: ~20 sec
  • SDSS g: VFLAT 4.0 + 20s x 10
    Execution Time: 11 min
  • SDSS u: QTH2 + 1s x 10
    Execution Time: 8 min

Total Execution time is about 20 min.

Prism Mode Calibrations

0945 - Yeah, OK, taking a break... It'sstill snowing, 100% humidity, -5C, wind 5-10m/s out of the SSW

Worked through lamp exposure times with the pinhole sieve, then switching to the LS5x60x0.6 segmented long slit

1035 - Prism Wavelength Calibration Lamps - Script: prlamps.cal

General Config: 3Kx3K ROI, Prisms, Dichroic, LS5x60x0.6 sllit mask

  • IMCS Center and Lock
    Execution Time: 45sec (gave it longer after the mode change even with presets)
  • Hg(Ar) + Ne Lamps x 3
    Blue: ND1.5 + 2s
    Red: ND1.5 + 1s
    Execution Time: 3 min
    Note: Sleep 30s to allow Hg lamp to warm up before start
  • Hg(Ar) + Ar Lamps x 3
    Blue: ND1.5 + 2s
    Red: ND1.5 + 1s
    Execution Time: 3 min
  • Kr Lamp x 3
    Blue: Clear + 5s
    Red: ND1.5 + 1s
    Execution Time: 3min
  • Xe Lamp x 3
    Blue: Clear + 20s
    Red: ND1.5 + 8s
    Execution Time: 4min
    Blue lags slightly due to longer exptime

Total execution time including IMCS and Hg(Ar) lamp warmup overheads: 14 minutes

1055 - Repeat of the prlamps.cal sequence to get the trace geometry (double-pass prisms have a small cross-dispersion term)

Execution time 14min as before.

1110 - Prism Spectral Flats - Script: prflats.cal

Pinhole Slit for the traces:

  • IMCS Center and Lock
    Execution Time: 45sec
  • QTH1 + ND1.5 x 10
    Blue: ND1.5 + 2s
    Red: Idle
    Execution Time: 8 m
  • QTH1 + UG5 x 10
    Blue: UG5 + 2s
    Red: Idle
    Execution Time: 8 m
  • VFLAT 4.0 + GG495 x 10
    Blue: Idle
    Red: GG495 + 3s
    Execution Time: 10m

Total Execution Time: 27min including IMCS overhead

1157 - LS5x60x0.6 Slit Mask, repeated prflats.cal flat sequence.

1201 - aborted sequence, reducing exptimes in blue to 1s, red to 1s because of saturation in the wider slit (oops)

UG5 is saturating in 1s in the peak in the 0.6-arcsec slit. May prove challenging to get blue flats, or we can find some way to work with saturated peaks? Also trying 0.5 sec exposure (minimum mechanical exptime) to see if that is consistent enough.

Looks like it is, but there is an odd behavior in the exposure countdown bar for exptimes < 1sec that will need investigating tomorrow (software bug I haven't seen before, perhaps because I've not used <1sec exposures in a long while).

Sequence completed without problems.

1240 - Ending Bias sequences

  • bias3k.cal
  • bias8K.cal

1300 - And done for the night.

-- RichardPogge - 16 Dec 2010
Topic revision: r6 - 18 Dec 2010, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback